What is the molar concentration of 4.63g of sugar C12H22O11 dissolved in 16.8 mL of solution?
1
Expert's answer
2019-01-10T04:04:24-0500

n = 4.63/342 = 0.0135 moles

C = 0.0135/0.0168 = 0.8M

Answer. 0.8M